# Duistarmaat’s triple index and the difference of Hermitian matrices

### v1

@@ 5. The Morse index of a difference of Hermitian matrices
In the case when $\ker A\subseteq\ker B$, the Hermitian matrices $A$, $B$, $A-B$ and $B^{+}-P_{B}A^{+}P_{B}$ are all identically zero on $\ker A$. This means we can restrict $A$ and $B$ (and thus $A-B$ and $B^{+}-P_{B}A^{+}P_{B}$) to the subspace $(\ker A)^{\perp}$, on which $A$ is invertible, without changing their Morse indices. It thus suffices to prove (5.5) when $A$ is invertible.
@@
To establish (5.6) under the assumption that $B$ is invertible, we use Theorem 1.4, equation (5.7), an analogue of (5.8) with $A$ and $B$ swapped, and, finally, (3.19). ∎
@@
Equation (1.1) with $-A$ in place of $A$ becomes
@@
Similar computations lead to a much simpler proof of [4, Prop. 6.1], which we give below.
@@
which establishes (1.1). Proposition 5.1 can be proved analogously, however, this requires an appropriate extension of the Haynsworth formula such as [15, Thm. 6.1] or [3, Thm. A.1].

### v2

@@ 5. The Morse index of a difference of Hermitian matrices
In the case when $\ker A\subseteq\ker B$, the Hermitian matrices $A$, $B$, ABA-B and $B^{+}-P_{B}A^{+}P_{B}$ are all identically zero on $\ker A$. This means we can restrict $A$ and $B$ (and thus ABA-B and $B^{+}-P_{B}A^{+}P_{B}$) to the subspace $(\ker A)^{\perp}$, on which $A$ is invertible, without changing their Morse indices. It thus suffices to prove (5.5) when $A$ is invertible.
@@
To establish (5.6) under the assumption that $B$ is invertible, we use Theorem 1.4, equation (5.7), an analogue of (5.8) with AA and $B$ swapped, and, finally, (3.19). ∎
@@
Equation (1.1) with A-A in place of AA becomes
@@
Similar computations lead to a much simpler proof of [5, Prop. 6.1], which we give below.
@@
which establishes (1.1). Proposition 5.1 can be proved analogously, however, this requires an appropriate extension of the Haynsworth formula such as [19, Thm. 6.1] or [4, Thm. A.1].